# Josephson effect in a Fibonacci quasicrystal

### v1

@@ Abstract
Quasiperiodicity has recently been proposed to enhance superconductivity and its proximity effect. At the same time, there has been significant experimental progress in the fabrication of quasiperiodic structures, also in reduced dimensions. Motivated by these developments, we use microscopic tight-binding theory to investigate the DC Josephson effect through a ballistic Fibonacci chain attached to two superconducting leads. The Fibonacci chain is one of the most studied examples of quasicrystals, hosting a rich multifractal spectrum, containing topological gaps with different winding numbers. We study how the Andreev bound states (ABS), current-phase relation, and the critical current depend on the quasiperiodic degrees of freedom, from short to long junctions. While the current-phase relation shows a traditional $2\pi$ sinusoidal or sawtooth profile, we find that the ABS obtain quasiperiodic oscillations and that the Andreev reflection is qualitatively altered, leading to quasiperiodic oscillations in the critical current as a function of junction length. Surprisingly, despite earlier proposals of enhanced superconductivity, we do not in general find an enhanced critical current. However, we find significant enhancement for reduced interface transparency due to the modified Andreev reflection. Furthermore, by varying the chemical potential, e.g. by an applied gate voltage, we find a fractal oscillation between superconductor-normal metal-superconductor (SNS) and superconductor-insulator-superconductor (SIS) behavior. Finally, we show that the winding of the subgap states leads to an equivalent winding in the critical current, such that the winding numbers, and thus the topological invariant, can be determined.
@@ I Introduction
Quasicrystals [1, 2, 3, 4, 5], and more generally aperiodic systems [6, 7, 8, 9], provide fascinating platforms to study exotic and topologically non-trivial behavior in physics [10, 11, 12, 13, 14, 15, 16, 17]. Quasicrystals are neither periodic nor random disordered, but instead quasiperiodic as they exhibiting a deterministic long-range order through discrete scale invariance and a non-crystallographic rotation symmetry [18, 19, 20]. In contrast to randomly disordered systems, quasicrystals therefore have clearly distinguished Bragg peaks in diffraction experiments [21, 22]. This peculiar combination leads to fascinating spectral and electronic properties [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], such as multifractality with critical states and wave functions that are neither localized nor extended [34, 35, 36], hyperuniformity [37, 38], pseudogaps [39, 40, 41, 42], topological gaps [43, 44, 45, 46], and topological invariants that are otherwise only possible in higher-dimensional systems [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61].
@@
The influence of quasiperiodicity on different ordered states and transport phenomena has also attracted significant interest [62, 50, 10, 11, 12, 13, 14, 15, 63, 17, 16, 64, 65]. For instance, it has recently been proposed that quasiperiodicity can enhance the superconducting order parameter and transition temperature [66, 67, 68, 69, 70], its proximity effect [71, 72], topological superconductivity [73], persistent currents in normal metal rings [74, 75], and cause enhanced or anomalous transport phenomena [76, 77, 78, 79, 80, 81]. The enhancement has been directly linked to the underlying topological [71] and critical states [66, 67, 68, 69] of the quasicrystal, similar to how criticality may enhance superconductivity also in disordered systems [82, 83, 84, 85, 86, 87, 88, 89]. However, while disordered systems typically exhibit such critical behavior only at the Anderson localization transition [90, 91], critical behavior is ubiquitous in quasicrystals [92].
@@
From an experimental standpoint, quasiperiodic systems are also becoming increasingly accessible. Significant progress in quasicrystal growth [93, 94] and synthetic engineering with atomic precision [95, 96, 97, 98, 99, 100, 101, 102, 103, 104] have recently enabled creation of quasiperiodic structures in reduced dimensions [105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117] and in moiré structures [118, 119, 120, 121]. A prime example is the Fibonacci quasicrystal [17], which is closely related to the dodecagonal and icosahedral quasicrystals [106], and hosts a multifractal spectrum of topological gaps with subgap winding states [59]. Its implementation as a one-dimensional (1D) atomic chain, the so-called Fibonacci chain, is also relevant to 3D systems [70] where it naturally appears or can be experimentally engineered in e.g. stacked materials [122, 123, 124, 125, 126]. Synthetic Fibonacci chains have additionally been realized in the context of photonics [127, 47, 49, 51, 128], phononics [129], polaritonics [130, 131, 132], cold atoms [133], dielectric chains [134], and magnonics [135]. In these systems, the Fibonacci chain spectrum and topology have often directly been measured [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61].
@@
Motivated by the experimental timeliness and the predictions of enhanced superconductivity, we here study the influence of quasiperiodicity on one of the most technologically important superconducting phenomena, namely the Josephson effect [136, 137, 138, 139]. In particular, we study a non-superconducting Fibonacci chain of length $L$ attached to two crystalline superconducting (S) leads using microscopic tight-binding theory, see Fig. 1(a). Since the Fibonacci chain is either normal conducting (N) or insulating (I), depending on if the Fermi level is outside or inside the topological gaps, respectively, we effectively study both ballistic SNS and SIS Josephson junctions. Furthermore, we study the short ($\xi_{0}\gg L$) to the long ($\xi_{0}\ll L$) junction regime, where $\xi_{0}$ is the superconducting coherence length [138]. We also model repeated Fibonacci chain supercells, i.e. so-called crystal approximants [21], see Figs. 1(b) and 1(c).
@@
We find that although the Josephson current-phase relation shows a conventional $2\pi$ sinusoidal or sawtooth profile, the low-energy Andreev bound states (ABS) are not conventional. Specifically, we demonstrate that the ABS obtain a quasiperiodic probability density, also at perfect resonance. Importantly, we also show how quasiperiodicity modifies the condition for zero-energy ABS, which generally generates the largest critical current [140], associated with the sawtooth current-phase profile and perfect Andreev reflection (i.e. zero normal reflection) [141]. We are in fact able to obtain a set of simple functional forms for this zero-energy condition, depending on the model parameters and spatial realization of the Fibonacci chain. As the junction length increases, one site at a time, the system changes quasiperiodically between these forms. Consequently, we find that the decay of the critical current with junction length is described by quasiperiodic oscillations, on top of the usual power law and exponential decays found in crystalline junctions [142].
@@
Surprisingly, while earlier studies have proposed that quasiperiodicity can enhance superconductivity and the proximity effect [66, 67, 68, 69, 70, 71, 72], we do not find that it generally enhances the critical current, at least not when compared to an idealized ballistic crystalline junction with a zero-energy state. However, beyond this idealized situation, we find that quasiperiodicity can cause a large enhancement of the critical current in junctions with poor transmission, due to it modifying the condition for zero-energy ABS. Moreover, we show that by applying a gate voltage, the junction varies between SNS and SIS behavior in a fractal manner, as the Fermi level enters and exits the topological gaps of the quasicrystal energy spectrum. Finally, we find that when the Fermi level is inside or close to the quasicrystal energy gaps, the winding of the subgap states leads to a similar winding in the critical current. We thus demonstrate how the Josephson current can probe the topological invariant in quasicrystals.
@@
The rest of this work is organized as follows. In Sec. II, we describe our model and relevant properties of the Fibonacci chain. In Sec. III we study the ABS spectrum and current-phase relation. In Sec. IV we study the critical current. In Sec. V we study the gate voltage dependence and also demonstrate how the critical current can measure the topological winding number. The work is concluded in Sec. VI.
@@ II.1 Fibonacci chain hopping model
In this subsection we define two methods to construct a Fibonacci chain, and our tight-binding model to study such a system, namely the Fibonacci hopping model [17].
@@
${H_{\mathrm{Q}}}=\sum_{\sigma,j\in\textrm{Q}}\left[\mu c_{j\sigma}^{\dagger}c_{j\sigma}-t_{j}(c_{j\sigma}^{\dagger}c_{j+1\sigma}+c^{\dagger}_{j+1\sigma}c_{j\sigma})\right],$ (3)
where $\mu$ is the chemical potential, i.e. controlled by the external gate voltage, $c^{\dagger}_{j\sigma}$ ($c_{j\sigma}$) is the creation (annihilation) operator of electronic states at site $j$ with spin $\sigma$, and the hopping integral $t_{j}$ takes one of the two values ${t_{\mathrm{A}}}$ or ${t_{\mathrm{B}}}$ following the Fibonacci chain, i.e. substituting the letters ${\rm A}\mapsto{t_{\mathrm{A}}}$ and ${\rm B}\mapsto{t_{\mathrm{B}}}$ following Eq. (2). From here on, ${t_{\mathrm{B}}}$ is our natural unit of energy. We introduce the hopping ratio $\rho\equiv{t_{\mathrm{A}}}/{t_{\mathrm{B}}}$ where $\rho\neq 1$ ($\rho=1$) corresponds to a quasiperiodic (crystalline) system. We note that the Fibonacci hopping model is closely related to the onsite Fibonacci model where an onsite potential instead varies quasiperiodically [23, 24, 17], with one main difference being that the Fibonacci number $F_{n}$ usually labels the number of bonds (sites) in the Fibonacci hopping (onsite) model.
@@ II.2 Fibonacci chain spectrum and topology
One of the most remarkable features of the Fibonacci chain is the opening (closing) of topological gaps at $\rho\neq 1$ ($\rho=1$), as illustrated in Fig. 2(a). We note that here $\mu=0$ corresponds to half-filling, see Fig. 2(b). Interestingly, at $\rho\neq 1$ there is a topological phase transition with the appearance of edge modes, see the red subgap state in Fig. 2(a), in analogy with the dimerized Su-Schrieffer-Heeger (SSH) model [144]. After all, a repeated Fibonacci approximant $C_{n}$ can in a sense be seen as a natural extension of the SSH model to include a richer sublattice structure, since the lowest-order non-trivial Fibonacci sublattice $C_{2}={t_{\mathrm{A}}}{t_{\mathrm{B}}}$ is similar to the the dimerized SSH chain. However, there are some important differences, like how the chain is usually terminated, and that the Fibonacci chain can host multiple topological gaps each with subgap states, Fig. 2(b). Furthermore, the exact subgap energy of the states depend directly on the phason angle as illustrated in Fig. 2(c). Additionally, the Fibonacci chain might have edge modes for both $\rho<1$ and $\rho>1$ in contrast to the SSH chain, and there are three (two) bands in the limit $\rho\to 0$ for the Fibonacci (SSH) chain, corresponding to the existence of three (two) kinds of nearest-neighbor hopping neighborhoods.
@@
The gap label $q$ was recently shown to be equivalent to a Chern number [47], which is easiest understood by first identifying the Fibonacci hopping model as topological equivalent with the 1D Aubry-André-Harper (AAH) model [145, 146, 147] and the 2D integer quantum Hall system [48, 52, 53]. Specifically, the Fibonacci chain has topological edge states which wind across the quasicrystal gaps labeled $q$, with corresponding winding number $q$. However, the states wind as a function of the phason angle $\phi$, not momentum as in crystalline systems. Specifically, as $\phi$ varies from 00 to $2\pi$, the topological states wind inside the gap $|q|$ times with direction $\operatorname{sgn}(q)$ as shown in Fig. 2(c). Due to the discrete nature of the characteristic function in Eq. (2), there are $F_{n}+1$ unique values (or phason flips) for the approximant $C_{n}$, which explains the step-like appearance of the energy levels in Fig. 2(c). Finally, we note that also the real-space localization of the winding state changes with $\phi$, such that the state winds back and forth across the chain $|q|$ times as shown in the probability density in Fig. 2(e). Recently, this was experimentally realized as topological pumping of the edge modes in phototonic [51] and polaritonic [132] systems. Finally, Fig. 2(f) shows a cut at fixed $\phi$, demonstrating that the spatial dependence of the edge modes is not typical exponential localization, but rather critical and furthermore with multifractal behavior [17]. The goal of our work is to see how these interesting properties influence the Josephson effect.
@@ II.3 Superconductivity and Josephson junction
The superconducting leads are modelled via the effective mean-field Hamiltonian
${H_{\mathrm{S}}}=$ $\sum_{\sigma,j\in\textrm{S}}\left[{\mu_{\mathrm{S}}}c^{\dagger}_{j\sigma}c_{j\sigma}-{t_{\mathrm{S}}}\left(c_{j\sigma}^{\dagger}c_{j+1\sigma}+c_{j+1\sigma}^{\dagger}c_{j\sigma}\right)\right]$
$-\sum_{j\in\textrm{S}}\left[\Delta_{j}c^{\dagger}_{j\uparrow}c^{\dagger}_{j\downarrow}+\Delta^{*}_{j}c_{j\downarrow}c_{j\uparrow}\right],$ (4)
with chemical potential ${\mu_{\mathrm{S}}}$ and hopping ${t_{\mathrm{S}}}$, and where $\uparrow$ ()($\downarrow$)( ↓ ) denotes spin up (down) states. We introduce the notation ${\rho_{\mathrm{S}}}\equiv{t_{\mathrm{S}}}/{t_{\mathrm{B}}}$, and for simplicity mainly focus on the scenario ${\mu_{\mathrm{S}}}=0$ (half-filling). We consider leads with ${N_{\mathrm{S}}}=135$ sites (we verify that increasing this size does not yield any noticeable difference on our results). We consider on-site $s$-wave superconductivity via the mean-field order parameter $\Delta_{j}$, given by the gap equation
$\Delta_{j}=U\sum_{\nu}u_{j\uparrow}^{n}v_{j\downarrow}^{n*}f(E_{\nu}),$ (5)
where $U=1.3{t_{\mathrm{B}}}$ is the attractive interaction yielding $|\Delta|\approx 0.06{t_{\mathrm{B}}}$ in the bulk of the SC leads at zero temperature, $\nu$ labels the eigenstates, $u$ and $v$ are the eigenvectors, and $f(E_{\nu})$ is the Fermi-Dirac distribution at energy level $E_{\nu}$. We have also carried out calculations with other $U$ up to $U=3{t_{\mathrm{B}}}$ and find qualitatively similar results.
@@
${H_{\mathrm{T}}}=-\sum_{\sigma,\langle i,j\rangle}{t_{\mathrm{int}}}c^{\dagger}_{j\sigma}c_{i\sigma}+{\rm H.c.},$ (6)
@@
where $\langle i,j\rangle$ denotes nearest neighbor sites $i\in\mathrm{Q}$ and $j\in\mathrm{S}$, and ${t_{\mathrm{int}}}$ is the interface hopping, as illustrated in Fig. 1(a). We investigate different values of ${t_{\mathrm{int}}}$, $\rho\equiv{t_{\mathrm{A}}}/{t_{\mathrm{B}}}$ and ${\rho_{\mathrm{S}}}\equiv{t_{\mathrm{S}}}/{t_{\mathrm{B}}}$ as a simplified way to model different interface transparencies, and note that this can still be related to results obtained with e.g. a scattering-matrix approach [141].
@@
We solve the resulting Hamiltonian $H={H_{\mathrm{Q}}}+{H_{\mathrm{S}}}+{H_{\mathrm{T}}}$ at zero temperature using the Bogoliubov-de Gennes method, for comparison both without and with self-consistency in the magnitude $|\Delta_{j}|$. Our criterion for self-consistency is that from self-consistency iteration number $m$ to ${m+i}$, the global relative error $\epsilon_{\mathrm{G}}=\left\|\Delta_{m+1}-\Delta_{m}\right\|_{2}/\left\|\Delta_{m}\right\|_{2}<10^{-7}$, where we use a Polyak convergence accelerator [148]. However, we generally find negligible influence of self-consistency apart from a small quantitative difference, which can be reproduced without self-consistency by simply adjusting the constant value $|\Delta_{j}|=\Delta_{0}$. We therefore present results for uniform bulk leads with $\Delta_{0}=0.06{t_{\mathrm{B}}}$. We fix the phases in the two superconducting leads using the phase difference $\Delta\theta\in[0,2\pi)$ as the relevant phase parameter. A finite phase difference $\Delta\theta>0$ leads to a supercurrent through the Josephson junction, which we compute via the bond charge current from site $k$ to $j$ via the current operator [149]
$I_{jk}=\frac{e}{i\hbar}\sum_{\nu,\sigma}\left[t_{jk}u_{j\sigma}^{n*}u_{k\sigma}^{n}f(E_{\nu})-t_{jk}^{*}u_{k\sigma}^{n*}u_{j\sigma}^{n}f(E_{\nu})\right],$ (7)
@@
with elementary charge $e$ and reduced Planck constant $\hbar$. The current contribution from individual energy levels $\nu$ are also obtained via the usual energy-phase dispersion at zero temperature $I_{\nu}(\Delta\theta)=-(2e/\hbar)(dE_{\nu}/d\Delta\theta)$ [150], and we verify that $I(\Delta\theta)=\sum_{\nu}I_{\nu}$ reproduces the same result as Eq. (7). The critical current ${I_{\mathrm{c}}}$ is defined as the maximum of the current-phase relation, ${I_{\mathrm{c}}}\equiv I(\Delta{\theta_{\mathrm{c}}})\equiv\max\{I(\Delta\theta):\Delta\theta\}$, where $\Delta{\theta_{\mathrm{c}}}$ is the critical phase difference.

### v2

@@ Abstract
Quasiperiodicity has recently been proposed to enhance superconductivity and its proximity effect. At the same time, there has been significant experimental progress in the fabrication of quasiperiodic structures, also in reduced dimensions. Motivated by these developments, we use microscopic tight-binding theory to investigate the DC Josephson effect through a ballistic Fibonacci chain attached to two superconducting leads. The Fibonacci chain is one of the most studied examples of quasicrystals, hosting a rich multifractal spectrum, containing topological gaps with different winding numbers. We study how the Andreev bound states (ABS), current-phase relation, and the critical current depend on the quasiperiodic degrees of freedom, from short to long junctions. While the current-phase relation shows a traditional $2\pi$ sinusoidal or sawtooth profile, we find that the ABS obtain quasiperiodic oscillations and that the Andreev reflection is qualitatively altered, leading to quasiperiodic oscillations in the critical current as a function of junction length. Surprisingly, despite earlier proposals of quasiperiodicity enhancing superconductivity compared to crystalline junctions, we do not in general find that it enhances the critical current. However, we find significant current enhancement for reduced interface transparency due to the modified Andreev reflection. Furthermore, by varying the chemical potential, e.g. by an applied gate voltage, we find a fractal oscillation between superconductor-normal metal-superconductor (SNS) and superconductor-insulator-superconductor (SIS) behavior. Finally, we show that the winding of the subgap states leads to an equivalent winding in the critical current, such that the winding numbers, and thus the topological invariant, can be determined.
@@ I Introduction
Quasicrystals [1, 2, 3, 4, 5], and more generally aperiodic systems [6, 7, 8, 9], provide fascinating platforms to study exotic and topologically non-trivial behavior in physics [10, 11, 12, 13, 14, 15, 16, 17, 18]. Quasicrystals are neither periodic nor random disordered, but instead quasiperiodic as they exhibit a deterministic long-range order through discrete scale invariance and a non-crystallographic rotation symmetry [19, 20, 21]. In contrast to randomly disordered systems, quasicrystals therefore have clearly distinguished Bragg peaks in diffraction experiments [22, 23]. This peculiar combination leads to fascinating spectral and electronic properties [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], such as multifractality with critical states and wave functions that are neither localized nor extended [35, 36, 37], hyperuniformity [38, 39], pseudogaps [40, 41, 42, 43], topological gaps [44, 45, 46, 47], and topological invariants that are otherwise only possible in higher-dimensional systems [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62].
@@
The influence of quasiperiodicity on different ordered states and transport phenomena has also attracted significant interest [63, 51, 10, 11, 12, 13, 14, 15, 64, 17, 16, 65, 66, 67, 68]. For instance, it has recently been proposed that quasiperiodicity can enhance the superconducting order parameter and transition temperature [69, 70, 71, 72, 73], its proximity effect [74, 75], topological superconductivity [76], persistent currents in normal metal rings [77, 78], and cause enhanced or anomalous transport phenomena [79, 80, 81, 82, 83, 84]. The enhancement has been directly linked to the underlying topological [74] and critical states [69, 70, 71, 72] of the quasicrystal, similar to how criticality may enhance superconductivity also in disordered systems [85, 86, 87, 88, 89, 90, 91, 92]. However, while disordered systems typically exhibit such critical behavior only at the Anderson localization transition [93, 94], critical behavior is ubiquitous in quasicrystals [95].
@@
From an experimental standpoint, quasiperiodic systems are also becoming increasingly accessible. Significant progress in quasicrystal growth [96, 97] and synthetic engineering with atomic precision [98, 99, 100, 101, 102, 103, 104, 105, 106, 107] have recently enabled creation of quasiperiodic structures in reduced dimensions [108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120] and in moiré structures [121, 122, 123, 124, 125]. A prime example is the Fibonacci quasicrystal [17], which is closely related to the dodecagonal and icosahedral quasicrystals [109], and hosts a multifractal spectrum of topological gaps with subgap winding states [60]. Its implementation as a one-dimensional (1D) atomic chain, the so-called Fibonacci chain, is also relevant to 3D systems [73] where it naturally appears or can be experimentally engineered in e.g. stacked materials [126, 127, 128, 129, 130]. Synthetic Fibonacci chains have additionally been realized in the context of photonics [131, 48, 50, 52, 132], phononics [133], polaritonics [134, 135, 136], cold atoms [137], dielectric chains or circuits [138, 139], and magnonics [140]. In these systems, the Fibonacci chain spectrum and topology have often directly been measured [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62].
@@
Motivated by the experimental timeliness and the predictions of enhanced superconductivity, we here study the influence of quasiperiodicity on one of the most technologically important superconducting phenomena, namely the Josephson effect [141, 142, 143, 144]. In particular, we study a non-superconducting Fibonacci chain of length $L$ attached to two crystalline superconducting (S) leads using microscopic tight-binding theory, see Fig. 1(a). Since the Fibonacci chain is either normal conducting (N) or insulating (I), depending on if the Fermi level is outside or inside the topological gaps, respectively, we effectively study both ballistic SNS and SIS Josephson junctions. Furthermore, we study the short ($\xi_{0}\gg L$) to the long ($\xi_{0}\ll L$) junction regime, where $\xi_{0}$ is the superconducting coherence length [143]. We also model repeated Fibonacci chain supercells, i.e. so-called crystal approximants [22], see Figs. 1(b) and 1(c). Overall, we perform extensive model calculations in these systems to address whether quasiperiodicity significantly influences the Josephson effect. Specifically, we systematically study how each model parameter influences the energy spectrum, current-phase relation, and critical current, contrasting the quasiperiodic and crystalline scenarios at each step.
@@
We find that although the Josephson current-phase relation shows a conventional $2\pi$ sinusoidal or sawtooth profile, the low-energy Andreev bound states (ABS) are not conventional. Specifically, we demonstrate that the ABS obtain a quasiperiodic probability density, also at perfect resonance. Importantly, we also show how quasiperiodicity modifies the condition for zero-energy ABS, which generally generates the largest critical current [145], associated with the sawtooth current-phase profile and perfect Andreev reflection (i.e. zero normal reflection) [146]. We are in fact able to obtain a set of simple functional forms for this zero-energy condition, depending on the model parameters and spatial realization of the Fibonacci chain. We find that the system changes quasiperiodically between these forms with junction length. Consequently, we find that the decay of the critical current with junction length is described by quasiperiodic oscillations, on top of the traditional power law and exponential decays found in crystalline junctions [147, 146, 148, 149]. These quasiperiodic oscillations imply significant sample-to-sample fluctuations, unless the junction is created with atomic precision, possible with modern STM techniques [117].
(b)${t_{\mathrm{A}}}$${t_{\mathrm{B}}}$${t_{\mathrm{A}}}$${t_{\mathrm{A}}}$${t_{\mathrm{B}}}$
@@
Surprisingly, while earlier studies have proposed that quasiperiodicity can enhance superconductivity and the proximity effect [69, 70, 71, 72, 73, 74, 75], we do not find that it generally enhances the critical current, at least not when compared to an idealized ballistic crystalline junction with a zero-energy state. However, beyond this idealized situation, we find that quasiperiodicity can cause a large enhancement of the critical current in junctions with poor transmission, due to it modifying the condition for zero-energy ABS. Moreover, we show that by applying a gate voltage, the junction varies between SNS and SIS behavior in a fractal manner, as the Fermi level enters and exits the topological gaps of the quasicrystal energy spectrum. Finally, we find that when the Fermi level is inside or close to the quasicrystal energy gaps, the topological subgap states can carry the majority of the Josephson current, while their winding leads to a similar winding in the critical current. We thus demonstrate how the Josephson current can probe the topological invariant in quasicrystals.
@@
The rest of this work is organized as follows. In Sec. II, we describe our model and relevant properties of the Fibonacci chain. In Sec. III we study how quasiperiodicity influences the ABS spectrum and current-phase relation. In Sec. IV we study the important critical current, and how it depends on the quasiperiodic modulation and interface transparency, the Fibonacci approximant order, and junction length. In Sec. V we study the gate voltage dependence and also demonstrate how the critical current can measure the topological winding number. The work is concluded in Sec. VI.
@@ II.1 Fibonacci chain hopping model
In this subsection we define how to construct a Fibonacci chain, and our tight-binding model to study such a system, namely the Fibonacci hopping model [17].
@@
${H_{\mathrm{Q}}}=\sum_{\sigma,j}\mu c_{j\sigma}^{\dagger}c_{j\sigma}-\sum_{\sigma,\langle ij\rangle}\left(t_{\langle ij\rangle}c_{j\sigma}^{\dagger}c_{i\sigma}+{\rm H.c.}\right),$ (3)
where $\mu$ is the chemical potential, e.g. controlled by an external gate voltage, $c^{\dagger}_{j\sigma}$ ($c_{j\sigma}$) is the creation (annihilation) operator of electronic states at site $j$ with spin $\sigma$, ${\langle ij\rangle}$ denotes nearest-neighbor sites $i$ and $j$ in the Fibonacci chain with hopping $t_{\langle ij\rangle}$ taking one of the two values ${t_{\mathrm{A}}}$ or ${t_{\mathrm{B}}}$ following the Fibonacci chain, i.e. substituting the letters ${\rm A}\mapsto{t_{\mathrm{A}}}$ and ${\rm B}\mapsto{t_{\mathrm{B}}}$ following Eq. (2). From here on, ${t_{\mathrm{B}}}$ is our natural unit of energy. We introduce the hopping ratio $\rho\equiv{t_{\mathrm{A}}}/{t_{\mathrm{B}}}$ where $\rho\neq 1$ ($\rho=1$) corresponds to a quasiperiodic (crystalline) system. We note that the Fibonacci hopping model is closely related to the onsite Fibonacci model where an onsite potential instead varies quasiperiodically [24, 25, 17], with one main difference being that the Fibonacci number $F_{n}$ usually labels the number of bonds (sites) in the Fibonacci hopping (onsite) model.
@@ II.2 Fibonacci chain spectrum and topology
One of the most remarkable features of the Fibonacci chain is the opening (closing) of topological gaps at $\rho\neq 1$ ($\rho=1$), as illustrated in Fig. 2(a). We note that here $\mu=0$ corresponds to half-filling, see Fig. 2(b). Interestingly, at $\rho\neq 1$ there is a topological phase transition with the appearance of edge modes, see the red subgap state in Fig. 2(a), in analogy with the dimerized Su-Schrieffer-Heeger (SSH) model [151]. After all, a repeated Fibonacci approximant $C_{n}$ can in a sense be seen as a natural extension of the SSH model to include a richer sublattice structure, since the lowest-order non-trivial Fibonacci sublattice $C_{2}={t_{\mathrm{A}}}{t_{\mathrm{B}}}$ is similar to the the dimerized SSH chain. However, there are some important differences, like how the chain is usually terminated, and that the Fibonacci chain can host multiple topological gaps each with subgap states, see Fig. 2(b). Furthermore, the exact subgap energy of the states depend directly on the phason angle as illustrated in Fig. 2(c). Additionally, the Fibonacci chain might have edge modes for both $\rho<1$ and $\rho>1$ in contrast to the SSH chain, and there are three (two) bands in the limit $\rho\to 0$ for the Fibonacci (SSH) chain, corresponding to the existence of three (two) kinds of nearest-neighbor hopping neighborhoods.
@@
The gap label $q$ was recently shown to be equivalent to a Chern number [48], which is easiest understood by first identifying the Fibonacci hopping model as topological equivalent with the 1D Aubry-André-Harper (AAH) model [152, 153, 154] and the 2D integer quantum Hall system [49, 53, 54]. Specifically, the Fibonacci chain has topological edge states which wind across the quasicrystal gaps labeled $q$, with corresponding winding number $q$. However, the states wind as a function of the phason angle $\phi$, not momentum as in crystalline systems. Specifically, as $\phi$ varies from 00 to $2\pi$, the topological states wind inside the gap $|q|$ times with direction $\operatorname{sgn}(q)$ as shown in Fig. 2(c). Due to the discrete nature of the characteristic function in Eq. (2), there are $F_{n}+1$ unique values (or phason flips) for the approximant $C_{n}$, which explains the step-like appearance of the energy levels in Fig. 2(c). Finally, we note that also the real-space localization of the winding state changes with $\phi$, such that the state winds back and forth across the chain $|q|$ times as shown in the probability density in Fig. 2(e). Recently, this was experimentally realized as topological pumping of the edge modes in phototonic [52] and polaritonic [136] systems. Finally, Fig. 2(f) shows a cut at fixed $\phi$, demonstrating that the spatial dependence of the edge modes is not extended or exponentially localized, but rather critical and furthermore with multifractal behavior [17]. The goal of our work is to see how the interesting quasiperiodic properties mentioned above influence the Josephson effect, by systematically investigating the model and parameter dependencies.
@@ II.3 Superconductivity and Josephson junction
Each superconducting lead is modelled via the effective mean-field Hamiltonian
${H_{\mathrm{S}}}=$ $\sum_{\sigma,j}{\mu_{\mathrm{S}}}c^{\dagger}_{j\sigma}c_{j\sigma}-\sum_{\sigma,\langle ij\rangle}\left({t_{\mathrm{S}}}c_{j\sigma}^{\dagger}c_{i\sigma}+{\rm H.c.}\right)$
$-\sum_{j}\left[\Delta_{j}c^{\dagger}_{j\uparrow}c^{\dagger}_{j\downarrow}+{\rm H.c.}\right],$ (4)
with chemical potential ${\mu_{\mathrm{S}}}$ and hopping ${t_{\mathrm{S}}}$, sites $i$ and $j$ in the respective superconducting lead, and where $\uparrow$ ()($\downarrow$)( ↓ ) denotes spin up (down) states. We introduce the notation ${\rho_{\mathrm{S}}}\equiv{t_{\mathrm{S}}}/{t_{\mathrm{B}}}$, and for simplicity mainly focus on the scenario ${\mu_{\mathrm{S}}}=0$ (half-filling). We consider superconducting leads with ${N_{\mathrm{S}}}=135$ sites (we verify that increasing this size does not yield any noticeable difference on our results), with on-site $s$-wave superconductivity via the mean-field order parameter $\Delta_{j}$.
@@
${H_{\mathrm{T}}}=-\sum_{\sigma,\langle ij\rangle}{t_{\mathrm{int}}}c^{\dagger}_{j\sigma}c_{i\sigma}+{\rm H.c.},$ (5)
@@
where ${t_{\mathrm{int}}}$ is the interface hopping between nearest neighbor sites ${\langle ij\rangle}$, one in the Fibonacci chain and the other in a superconducting lead as illustrated in Fig. 1(a). We investigate different values of ${t_{\mathrm{int}}}$, $\rho\equiv{t_{\mathrm{A}}}/{t_{\mathrm{B}}}$ and ${\rho_{\mathrm{S}}}\equiv{t_{\mathrm{S}}}/{t_{\mathrm{B}}}$ as a simplified way to model different interface transparencies, and note that this can still be related to results obtained with e.g. a scattering-matrix approach [146].
@@
We solve the resulting Hamiltonian $H={H_{\mathrm{Q}}}+{H_{\mathrm{S}}}+{H_{\mathrm{T}}}$ at zero temperature using the Bogoliubov-de Gennes method, using uniform superconductivity in the leads $|\Delta_{j}|=\Delta_{0}=0.06{t_{\mathrm{B}}}$. We verify that self-consistency, capturing the inverse proximity effect, does not qualitatively modify the results  [^1] We diagonalize the Hamiltonian both with and without self-consistency using a homogeneous bulk magnitude $|\Delta_{j}|$. Our criterion for self-consistency is that from self-consistency iteration number $m$ to ${m+i}$, the global relative error $\epsilon_{\mathrm{G}}=\left\|\Delta_{m+1}-\Delta_{m}\right\|_{2}/\left\|\Delta_{m}\right\|_{2}<10^{-7}$, where we use a Polyak convergence accelerator [167]. We generally find negligible influence of self-consistency apart from a small quantitative difference, which can be reproduced without self-consistency by simply adjusting $\Delta_{0}$.. We fix the phases in the two superconducting leads using the phase difference $\Delta\theta\in[0,2\pi)$ as the relevant phase parameter. A finite phase difference $\Delta\theta>0$ leads to a supercurrent through the Josephson junction, which we compute via the bond charge current from site $k$ to $j$ via the current operator [156]
$I_{jk}=\frac{e}{i\hbar}\sum_{\nu,\sigma}\left[t_{jk}u_{j\sigma}^{\nu*}u_{k\sigma}^{\nu}f(E_{\nu})-t_{jk}^{*}u_{k\sigma}^{\nu*}u_{j\sigma}^{\nu}f(E_{\nu})\right],$ (6)
with elementary charge $e$, reduced Planck constant $\hbar$, eigenvectors $u$ and $v$, Fermi-Dirac distribution $f(E_{\nu})$, and $\nu$ labels the eigenstates.
@@
The current contribution from individual energy levels $\nu$ are also obtained via the usual energy-phase dispersion at zero temperature $I_{\nu}(\Delta\theta)=-(2e/\hbar)(dE_{\nu}/d\Delta\theta)$ [157], and we verify that $I(\Delta\theta)=\sum_{\nu}I_{\nu}$ reproduces the same result as Eq. (6). The critical current ${I_{\mathrm{c}}}$ is defined as the maximum of the current-phase relation,
${I_{\mathrm{c}}}\equiv I(\Delta{\theta_{\mathrm{c}}})\equiv\max\{I(\Delta\theta):\Delta\theta\},$ (7)
where $\Delta{\theta_{\mathrm{c}}}$ is the critical phase difference.